OpenStudy (anonymous):

Really need help please!!!!! Find an equivalent function to f(x) = 3(6)^2^x.

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Please help me

OpenStudy (irishboy123):

\(\large f(x) = 3(6)^{2^x} \) is it what you meant?

OpenStudy (anonymous):

well its to the power of 2x @IrishBoy123

OpenStudy (irishboy123):

yet you posted: 3(6)^2***^x*** are you sure? do you have a screengrab?

OpenStudy (irishboy123):

but you now say: \(\huge f(x) = 3(6)^{2x}\)

OpenStudy (anonymous):

sry

OpenStudy (irishboy123):

no worries by equivalent function, do you have any idea what that means? you can spin this any old way but is there a thing you are learning that we need to keep in mind?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

f(x) = 3(36)x f(x) = 182x f(x) = 108x f(x) = 9x(36x) @IrishBoy123

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

so it's multiple choice. did you mention that?

OpenStudy (irishboy123):

\(\large f(x) = 3(6)^{2x}\) some basic rules: \(\large \{ \#1\} \ \ a^m . a^n = a^{m + n}\) \( \large \{ \#2\} \ \ (a^{m})^n = (a^{n})^m = a^{m \times n} \) so \( 3 \ (6)^{2x} = 3 \ (6^2)^x = 3 \ (6^x)^2\) and \(6^2 = 36\) that is going to be all you need